A BALL IS PROJECTED VERTICALLY UPWARDS . ITS POTENTIAL ENERGY AND KINETIC ENERGY AT A POINT ARE IN THE RATIO 4: 3. IF THE SAME BALL IS PROJECTED WITH DOUBLE  THE SPEED THEN THE RATIO OF KINETIC AND POTENTIAL ENERGY IS WHAT AT THE SAME POINT ?

Grade:12

1 Answers

Eshan
5 years ago
Let the kinetic energy beK=\dfrac{1}{2}mv^2
Hence potential energy=P=\dfrac{4}{3}K

If the speed of ball is doubled, kinetic energy becomesK'=\dfrac{1}{2}m(2v)^2=4K
But since the height of the ball in the second case is same, the potential energy of the ball is same.

Hence the new ratio=\dfrac{K'}{P}=\dfrac{4K}{4K/3}=3:1